Which of the following is true of line​ management?

A.

It is directly responsible for achieving the goals of the organization.

B.

It provides​ advice, support, and assistance to staff management.

C.

It is responsible of management accounting functions.

D.

It only includes the top level management.

Cost accumulation and cost assignment are synonymous.

True

False

Which of the following is an assumption of CVP​ analysis?

A.The unit−selling price is variable as it is subject to demand and supply.

B.When​ graphed, total costs curve upward.

C.Total costs can be divided into inventoriable and period costs with respect to the level of output.

D.Total costs can be divided into a fixed component and a component that is variable with respect to the level of output.

The East Company manufactures several different products. Unit costs associated with Product ORD105 are as​ follows:

Direct materials

​$54

Direct manufacturing labor

8

Variable manufacturing overhead

11

Fixed manufacturing overhead

25

Sales commissions​ (2% of​ sales)

5

Administrative salaries

12

     Total

​$115

What is the percentage of the total fixed costs per unit associated with Product ORD105 with respect to total​ cost?

A.​26%

B.​32%

C.​37%

D.​15%

If a company is planning to reduce the selling price per unit but not reduce expectations of​ revenues, they must believe that​ ________.

A.increased collections will increase income

B.variable costs will decline as well

C.more units will be sold

D.the fixed costs will cover the lower sales price
Answer #1

A. Answer is (a). They are directly responsible of achievning organisation's goals. A line manager is an employee who directly manages other employees and operations. Therefor, they dont handle accounting functions but are involved in production process. They are also not the top level managers.

B. False. ost accumulation is the process of linking costs with the units produced. cost assignment decides whether costs must be assigned to jobs or processes.

C. Answer is (D).Total costs can be divided into a fixed component and a component that is variable with respect to the level of output.

D. Fixed cost= Fixed manufacturing overhead+administrative salaries= 25+12= $37. Total cost= $115. Percentage of fixed cost to total cost= 37/115= 32%. Answer is (b)